OpenStudy (aaronandyson):

A bag has 24 balls.sme are white and the remaining are black.if the probability of drawing a white balck is twice that of a black ball. find the no. of white balls.

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

Number of white balls: \(x\). Number of black balls: \(24-x\). P(white) = \(\frac{x}{24}\) P(black) = \(\frac{24-x}{24}\) Now use the piece of information you're given and you'll get a very simple equation.

OpenStudy (aaronandyson):

karna kya hai...?

Parth (parthkohli):

"The probability of finding a white ball is twice that of finding a black ball." P(white) = x/24 P(black) = (24-x)/24 P(white) = 2*P(black) => x/24 = 2*(24-x)/24

OpenStudy (aaronandyson):

lemme solve

OpenStudy (aaronandyson):

x/24 = 24-x/12

OpenStudy (aaronandyson):

No. of white balls = 16

Parth (parthkohli):

Yes, correct.

OpenStudy (aaronandyson):

Thanks man.

OpenStudy (aaronandyson):

fan karo muhje aapse zaruri baat karni hai